# PyMieDiff: A differentiable Mie scattering library

### v1

@@ Abstract
Light scattering by spherical-shaped particles of sizes comparable to the wavelength is foundational in many areas of science, from chemistry to atmospheric science, photonics and nanotechnology. With the new capabilities offered by machine learning, there is a great interest in end-to-end differentiable frameworks for scattering calculations. Here we introduce PyMieDiff, a fully differentiable, GPU‑compatible implementation of Mie scattering for core‑shell particles in PyTorch. The library provides native, autograd‑compatible spherical Bessel and Hankel functions, vectorized evaluation of Mie coefficients, and APIs for computing efficiencies, angular scattering, and near‑fields. All inputs – geometry, material dispersion, wavelengths, and observation angles and positions – are represented as tensors, enabling seamless integration with gradient‑based optimisation or physics‑informed neural networks. The toolkit can also be combined with “TorchGDM” for end-to-end differentiable multi-particle scattering simulations. PyMieDiff is available under an open source licence at https://github.com/UoS-Integrated-Nanophotonics-group/MieDiff.
@@ I Introduction
Mie theory provides an analytical solution for light scattering by spherical particles, which underpins many nano-optical applications (e.g., color generation, dielectric metamaterials, nanoantennas, radiative cooling).mieBeitrageZurOptik1908 Multi-layer (core–shell) spheres extend this to complex nanoparticle designs. While the forward scattering problem can be solved by standard Mie formulae,bohrenAbsorptionScatteringLight1998 the inverse design of such particles remains challenging: it typically requires many repeated forward simulations and costly optimization. Recent work has therefore turned to machine learning, training neural nets to predict scattering spectra and act as differentiable surrogates.kuhnInverseDesignCoreshell2022 ; peurifoyNanophotonicParticleSimulation2018 ; soSimultaneousInverseDesign2019 ; estrada-realInverseDesignFlexible2022 ; sounGradientbasedOptimizationCoreshell2025 An alternative approach, which involves integrating the exact Mie solution directly into gradient-based design workflows has not been fully exploited, as analytic derivatives tend to become very bulky,liComputationMieDerivatives2013 especially for core-shell spheres.
@@
A variety of open-source packages implement Mie theory for spheres. For example, “MiePython” is a pure-Python (NumPy/Numba) implementation of Mie theory for homogeneous spheres.prahlMiepythonPythonLibrary2025 “PyMieSim” is an open-source Python/C++ toolkit that supports scattering from spheres, infinite cylinders and core–shell geometries.sivry-houlePyMieSimOpensourceLibrary2023 “Scattnlay” is a c++ programm, focused on the calculation of nearfields inside and around multi-layer spherical particles.ladutenkoMieCalculationElectromagnetic2017 “pyMieCS” is a vectorized NumPy library specialized for core–shell nanoparticles.wiechaPymiecsSimplePython2024 These tools offer validated, high-speed forward solvers, but none provide native differentiation capabilities or GPU backends.
@@
Automatic differentiation (AD) is the key numerical technique behind deep learning, and allows to calculate arbitrary derivatives for any numerical calculation with close to analytical precision.wengertSimpleAutomaticDerivative1964 The photonics community has seen rapid growth of differentiable simulation tools based on AD. Recent efforts have produced open-source Maxwell solvers with autodifferentiation (e.g., FEM, Fourier modal, FDTD or volume integral approaches) to enable gradient-based inverse design and topology optimization.hughesForwardModeDifferentiationMaxwells2019 ; minkovInverseDesignPhotonic2020 ; colburnInverseDesignFlexible2021 ; wangDifferentiableEngineDeep2022 ; luceMergingAutomaticDifferentiation2024 ; kimMeentDifferentiableElectromagnetic2024 ; ponomarevaTorchGDMGPUAcceleratedPython2025 For Mie computations however, so far still external libraries are required, which are not AD-compatible.ponomarevaTorchGDMGPUAcceleratedPython2025
@@
Recent works have highlighted the power of integrating physics-based models with deep learning. An important example are deep learning based nanoparticle design approaches which, so far, often rely on data-based tandem networks or surrogate models.liuTrainingDeepNeural2018 ; soSimultaneousInverseDesign2019 ; estrada-realInverseDesignFlexible2022 ; khaireh-waliehNewcomersGuideDeep2023 These solutions often suffer from a lack of fidely of the surrogate models dinsdaleDeepLearningEnabled2021 , as their interpolation is limited to the dataset they are trained on. These large datasets only cover a limited parameter space, and new problems reqiure new datasets. Replacing such surrogate models by AD capable analytical solvers would suppress this typical point of failure.
@@
We present the PyTorch-based toolkit “PyMieDiff” which implements analytical Mie theory for core–shell spherical particles with full support for automatic differentiation. Our library implements the standard Mie recurrence (spherical Bessel/Hankel forward and backward recurrences, angular functions, vector spherical harmonicscaiComputationSphericalBessel2011 ) entirely in PyTorch. All quantities – particle geometry (layer thicknesses, materials), wavelengths, and scattering angles – can be treated as differentiable tensors. As a result any Mie-derived observable can be backpropagated to compute gradients with respect to any input parameter. The code is fully vectorized over Mie orders, wavelengths, angles, and positions. It supports batched evaluation of many particles in parallel and runs on GPU. Our simple API exposes a Particle class for defining core-shell spheres with possibly dispersive materials (using the refractiveindex.info formatpolyanskiyRefractiveindexinfoDatabaseOptical2024 ). It implements methods to compute extinction, absorption and scattering efficiencies, far-fields and S-matrix elements, as well as near-fields, all with autograd support, as illustrated in figure 1.
@@
  • Compatibility with TorchGDM for multi-particle scattering simulations with autodiff support.ponomarevaTorchGDMGPUAcceleratedPython2025
@@
By embedding analytic Mie theory in an autodiff framework, our toolkit enables new design approaches in nano-optics. For example, we demonstrate using the analytic forward model in gradient-based inverse design of core–shell particles (including a “tandem” neural network architecture whose forward layer is our Mie solver). Because the Mie calculations are exact, this avoids the need for approximate surrogate training.estrada-realInverseDesignFlexible2022 It also opens the door to differentiable sensitivity analyses and rapid parameter retrieval from data.
@@
  • Machine learning integration: embedding the Mie forward model as a layer in neural networks (e.g., tandem or physics-informed networks) so that analytic gradients flow through the scattering physics for training.jiangMultiobjectiveCategoricalGlobal2020
@@ II Implementation
To obtain $j_{n}$, we run through Eq. (3) as backward recurrence, starting from seeds $j_{n}=0$ and $j_{n+1}=10^{-25}$ at a sufficiently large ordercaiComputationSphericalBessel2011 . The sequence is finally re-normalized to the known $j_{0}$ ($j_{0}(z)=\sin z/z$). The Neumann functions $y_{n}$ are generated via the same three-term recurrence, applied upward from analytic seeds ($y_{0}(z)=-\cos z/z$, $y_{1}(z)=-\cos z/z^{2}-\sin z/z$).gillmanAccurateRecursiveGeneration1988 Implementing both branches using only differentiable PyTorch tensor operations preserves automatic differentiation, while avoiding overflow/underflow and the loss of accuracy encountered with lossy materials, for example in plasmonic parameter regimes. Spherical Hankel functions and derivatives are generated from the spherical Bessel relations (see Eqs. (2) and appendix .4).
@@
Figure 3: Mie forward evaluation by the example of scattering from a gold-silicon core-shell particle with core radius $r_{c}=20\,$nm and shell radius $r_{s}=100\,$nm, placed in vacuum and illuminated by a linearly polarized plane wave. Tabulated material permittivities are taken from literature.johnsonOpticalConstantsNoble1972 ; edwardsSiliconSi1997 (a) Extinction (blue), scattering (orange), and absorption (green) efficiency spectra. (b) Scattering radiation patterns in the scattering plane at selected wavelengths for perpendicularly (blue lines) and parallel (orange dashed lines) polarized light. (c) Near-fields (electric and magnetic in top and bottom row) inside and around the particle, evaluated at the resonance at $\lambda_{0}=575\,$nm. The shown maps are $400\times 400\,$nm2. Colorbars indicate field intensity (leftmost panels) and amplitude (second from left to right panels), relative to the incident field absolute amplitude.
@@ III Benchmark
On multi-core CPUs, PyMieDiff demonstrates good parallel efficiency, which seems to approach a plateau after eight cores (Fig. 2c). These parallelization capabilities are automatically used since they are inherent to PyTorch. This scaling behaviour is particularly beneficial for large-scale parameter sweeps, many-particle optimization tasks or machine learning integration with batched training schemes. In a direct comparison with other publicly available toolkits (Fig. 2d), PyMieDiff achieves markedly lower evaluation times per wavelength as soon as batched evaluation is possible, when our toolkit provides more than one order-of-magnitude improvement over the other evaluated tools. Note that while the T-Matrix package “treams”beutelTreamsTmatrixbasedScattering2024 is included for completeness, it is not parallelized and not specifically optimized for single particle Mie calculations.
@@ IV.2 Gradient based optimization
In this example, the popular Adam optimizer is used.kingmaAdamMethodStochastic2014 While Adam is ideal for mini-batch based optimization as typically used for neural network training, it performs well also on our optimization task. We demonstrate in the online documentation how to use the L-BFGS optimizer as an alternative.nocedalUpdatingQuasiNewtonMatrices1980
@@ IV.3 Core-shell design: Mie informed tandem model
As a simple technical example, we train a design network, capable to predict core-shell geometries that fulfill given target optical properties. This so-called “tandem” model is a commonly used deep learning method for inverse design.liuTrainingDeepNeural2018 ; khaireh-waliehNewcomersGuideDeep2023 It works by regularizing the training of the ill-posed inverse problem through a forward predictor, as depicted in figure 5a. Typically, the forward model is a trained neural network that approximates a phyics solver, adding automatic differentiation capabilities. With PyMieDiff as fast, autograd-compatible Mie solver, analytical Mie theory can be used as an error-free forward model through which the loss can be backpropagated to train the inverse network.
@@ IV.4 Autodiff multi-scattering: End-to-end metasurface design
As a final example, we demonstrate how PyMieDiff can be combined with the autodiff light scattering simulation toolkit “TorchGDM”,ponomarevaTorchGDMGPUAcceleratedPython2025 to perform multi-particle scattering simulations. PyMieDiff evaluates the Mie response of one or several core-shell particles, which are then used by torchGDM to generate a structure model and calculate the optical response of an ensemble of many scatterers, where all optical interactions between the particles are taken into account. Since both tools fully support PyTorch automatic differentiation, gradient optimization can be used to iteratively optimize the positions and core/shell size parameters for any target observable. This is illustrated in figure 6a.
@@ V Limitations and perspectives
  • The recurrences currently require to be performed at double precision, which significantly slows down the code on consumer grade GPUs. In the future, more stable algorithms may be implemented such as logarithmic derivatives.majicNumericallyStableFormulation2020 ; zhangModifiedRecursiveTransfer2024
  • Vector spherical harmonics are currently implemented only for order $l=1$, which limits the field calculations to particles with spherical symmetry. We plan to add pure pytorch implementations of general vector spherical harmonics that could be used for field calculations of general T-matrices. A possible route could be to expand on NVIDIA’s recent work about spherical Fourier neural operators.bonevSphericalFourierNeural2023
  • The current recurrences can become unstable in cases where many expansion orders are contributing to the response, for instance with very large particles or plasmonic/dielectric interfaces in larger particles. Be aware of these limitations. This could be mitigated in the future by implementing more stable algorithms, such as logarithmic derivatives recurrences.bohrenAbsorptionScatteringLight1998
@@ VI Conclusions
At the time of the submission, we became aware of a manuscript developing similar ideas.dreams_paper_to_be_added This underpins the timeliness and importance of the differentiable formulation of algorithms that solve multiple-scattering problems, a key requirement for inverse design of photonic nanostructures.
@@ Acknowledgements.
For comparison and testing, we provide an alternative PyTorch-based scheme to compute the spherical Bessel functions $j_{n}$ using a continued fraction calculation,lentzContinuedFractionCalculation1990 Using the Riccati–Bessel form $\psi_{n}(z)=z\,j_{n}(z)$ the three-term recurrence reads
@@
We truncate the continued fraction at a sufficiently large $n_{\max}$,caiComputationSphericalBessel2011 compute the $A_{n}$ by backward recurrence, reconstruct $\psi_{n}$ (and hence $j_{n}=\psi_{n}/z$) up to a multiplicative constant, and fix the global scale using $j_{0}(z)=\sin z/z$. Both implementations for $j_{n}$ are similarly stable and can be selected with the argument which_jn (“ratios” or “recurrence”).
@@
The vector spherical harmonics $\mathbf{N}_{lm}^{(k)}$ and $\mathbf{M}_{lm}^{(k)}$ can be used to describe the near fields of Mie scattering.asadovaTmatrixRepresentationOptical2025 For spherical symmetry, we need only the $l=1$ odd and even order harmonics, which are given by:bohrenAbsorptionScatteringLight1998
@@
The Mie coefficients for a core-shell spherical particle are given bybohrenAbsorptionScatteringLight1998
@@
The Mie coefficients for internal fields can be found in literature.bohrenAbsorptionScatteringLight1998
@@
where $\mu=\cos\theta$. The angular functions $\pi_{n}(\cos\theta)$ and $\tau_{n}(\cos\theta)$ are defined from the associated Legendre polynomials $P_{n}^{1}(\cos\theta)$.bohrenAbsorptionScatteringLight1998 These satisfy the recurrence relations
@@
The near-fields can be found by matching the continuity conditions at each spherical interface, which yields following formulae:bohrenAbsorptionScatteringLight1998

### v2

@@ Abstract
Light scattering by spherical-shaped particles of sizes comparable to the wavelength is foundational in many areas of science, from chemistry to atmospheric science, photonics and nanotechnology. With the new capabilities offered by machine learning, there is a great interest in end-to-end differentiable frameworks for scattering calculations. Here we introduce PyMieDiff, a fully differentiable, GPU-compatible implementation of Mie scattering for core-shell particles in PyTorch. The library provides native, autograd-compatible spherical Bessel and Hankel functions, vectorized evaluation of Mie coefficients, and APIs for computing efficiencies, angular scattering, and near-fields. All inputs – geometry, material dispersion, wavelengths, and observation angles and positions – are represented as tensors, enabling seamless integration with gradient-based optimisation or physics-informed neural networks. The toolkit can also be combined with “TorchGDM” for end-to-end differentiable multi-particle scattering simulations. PyMieDiff is available under an open source licence at https://github.com/UoS-Integrated-Nanophotonics-group/MieDiff.
@@ I Introduction
Mie theory provides an analytical solution for light scattering by spherical particles, which underpins many nano-optical applications (e.g., color generation, dielectric metamaterials, nanoantennas, radiative cooling).Mie (1908) Multi-layer (core–shell) spheres extend this to complex nanoparticle designs. While the forward scattering problem can be solved by standard Mie formulae,Bohren and Huffman (1998) the inverse design of such particles remains challenging: it typically requires many repeated forward simulations and costly optimization. Recent work has therefore turned to machine learning, training neural nets to predict scattering spectra and act as differentiable surrogates.Kuhn et al. (2022); Peurifoy et al. (2018); So et al. (2019); Estrada-Real et al. (2022); Soun et al. (2025) An alternative approach, which involves integrating the exact Mie solution directly into gradient-based design workflows has not been fully exploited, as analytic derivatives tend to become very bulky,Li and Bowler (2013) especially for core-shell spheres.
@@
A variety of open-source packages implement Mie theory for spheres. For example, “MiePython” is a pure-Python (NumPy/Numba) implementation of Mie theory for homogeneous spheres.Prahl (2025) “PyMieSim” is an open-source Python/C++ toolkit that supports scattering from spheres, infinite cylinders and core–shell geometries.de Sivry-Houle et al. (2023) “Scattnlay” is a c++ programm, focused on the calculation of nearfields inside and around multi-layer spherical particles.Ladutenko et al. (2017) “pyMieCS” is a vectorized NumPy library specialized for core–shell nanoparticles.Wiecha (2024) These tools offer validated, high-speed forward solvers, but none provide native differentiation capabilities or GPU backends.
@@
Automatic differentiation (AD) is the key numerical technique behind deep learning, and allows to calculate arbitrary derivatives for any numerical calculation with close to analytical precision.Wengert (1964) The photonics community has seen rapid growth of differentiable simulation tools based on AD. Recent efforts have produced open-source Maxwell solvers with autodifferentiation (e.g., FEM, Fourier modal, FDTD or volume integral approaches) to enable gradient-based inverse design and topology optimization.Hughes et al. (2019); Minkov et al. (2020); Colburn and Majumdar (2021); Wang et al. (2022); Luce et al. (2024); Kim et al. (2024); Ponomareva et al. (2025) For Mie computations however, so far still external libraries are required, which are not AD-compatible.Ponomareva et al. (2025)
@@
Recent works have highlighted the power of integrating physics-based models with deep learning. An important example are deep learning based nanoparticle design approaches which, so far, often rely on data-based tandem networks or surrogate models.Liu et al. (2018); So et al. (2019); Estrada-Real et al. (2022); Khaireh-Walieh et al. (2023) These solutions often suffer from a lack of fidely of the surrogate models Dinsdale et al. (2021), as their interpolation is limited to the dataset they are trained on. These large datasets only cover a limited parameter space, and new problems reqiure new datasets. Replacing such surrogate models by AD capable analytical solvers would suppress this typical point of failure.
@@
We present the PyTorch-based toolkit “PyMieDiff” which implements analytical Mie theory for core–shell spherical particles with full support for automatic differentiation. Our library implements the standard Mie recurrence (spherical Bessel/Hankel forward and backward recurrences, angular functions, vector spherical harmonicsCai (2011)) entirely in PyTorch. All quantities – particle geometry (layer thicknesses, materials), wavelengths, and scattering angles – can be treated as differentiable tensors. As a result any Mie-derived observable can be backpropagated to compute gradients with respect to any input parameter. The code is fully vectorized over Mie orders, wavelengths, angles, and positions. It supports batched evaluation of many particles in parallel and runs on GPU. Our simple API exposes a Particle class for defining core-shell spheres with possibly dispersive materials (using the refractiveindex.info formatPolyanskiy (2024)). It implements methods to compute extinction, absorption and scattering efficiencies, far-fields and S-matrix elements, as well as near-fields, all with autograd support, as illustrated in figure 1.
@@
  • Compatibility with TorchGDM for multi-particle scattering simulations with autodiff support.Ponomareva et al. (2025)
@@
By embedding analytic Mie theory in an autodiff framework, our toolkit enables new design approaches in nano-optics. For example, we demonstrate using the analytic forward model in gradient-based inverse design of core–shell particles (including a “tandem” neural network architecture whose forward layer is our Mie solver). Because the Mie calculations are exact, this avoids the need for approximate surrogate training.Estrada-Real et al. (2022) It also opens the door to differentiable sensitivity analyses and rapid parameter retrieval from data.
@@
  • Machine learning integration: embedding the Mie forward model as a layer in neural networks (e.g., tandem or physics-informed networks) so that analytic gradients flow through the scattering physics for training.Jiang and Fan (2020)
@@ II Implementation
To obtain $j_{n}$, we run through Eq. (3) as backward recurrence, starting from seeds $j_{n}=0$ and $j_{n+1}=10^{-25}$ at a sufficiently large orderCai (2011). The sequence is finally re-normalized to the known $j_{0}$ ($j_{0}(z)=\sin z/z$). The Neumann functions $y_{n}$ are generated via the same three-term recurrence, applied upward from analytic seeds ($y_{0}(z)=-\cos z/z$, $y_{1}(z)=-\cos z/z^{2}-\sin z/z$).Gillman and Fiebig (1988) Implementing both branches using only differentiable PyTorch tensor operations preserves automatic differentiation, while avoiding overflow/underflow and the loss of accuracy encountered with lossy materials, for example in plasmonic parameter regimes. Spherical Hankel functions and derivatives are generated from the spherical Bessel relations (see Eqs. (2) and appendix .4).
@@
Figure 3: Mie forward evaluation by the example of scattering from a gold-silicon core-shell particle with core radius $r_{c}=20\,$nm and shell radius $r_{s}=100\,$nm, placed in vacuum and illuminated by a linearly polarized plane wave. Tabulated material permittivities are taken from literature.Johnson and Christy (1972); Edwards (1997) (a) Extinction (blue), scattering (orange), and absorption (green) efficiency spectra. (b) Scattering radiation patterns in the scattering plane at selected wavelengths for perpendicularly (blue lines) and parallel (orange dashed lines) polarized light. (c) Near-fields (electric and magnetic in top and bottom row) inside and around the particle, evaluated at the resonance at $\lambda_{0}=575\,$nm. The shown maps are $400\times 400\,$nm2. Colorbars indicate field intensity (leftmost panels) and amplitude (second from left to right panels), relative to the incident field absolute amplitude.
@@ III Benchmark
On multi-core CPUs, PyMieDiff demonstrates good parallel efficiency, which seems to approach a plateau after eight cores (Fig. 2c). These parallelization capabilities are automatically used since they are inherent to PyTorch. This scaling behaviour is particularly beneficial for large-scale parameter sweeps, many-particle optimization tasks or machine learning integration with batched training schemes. In a direct comparison with other publicly available toolkits (Fig. 2d), PyMieDiff achieves markedly lower evaluation times per wavelength as soon as batched evaluation is possible, when our toolkit provides more than one order-of-magnitude improvement over the other evaluated tools. Note that while the T-Matrix package “treams”Beutel et al. (2024) is included for completeness, it is not parallelized and not specifically optimized for single particle Mie calculations.
@@ IV.2 Gradient based optimization
In this example, the popular Adam optimizer is used.Kingma and Ba (2014) While Adam is ideal for mini-batch based optimization as typically used for neural network training, it performs well also on our optimization task. We demonstrate in the online documentation how to use the L-BFGS optimizer as an alternative.Nocedal (1980)
@@ IV.3 Core-shell design: Mie informed tandem model
As a simple technical example, we train a design network, capable to predict core-shell geometries that fulfill given target optical properties. This so-called “tandem” model is a commonly used deep learning method for inverse design.Liu et al. (2018); Khaireh-Walieh et al. (2023) It works by regularizing the training of the ill-posed inverse problem through a forward predictor, as depicted in figure 5a. Typically, the forward model is a trained neural network that approximates a phyics solver, adding automatic differentiation capabilities. With PyMieDiff as fast, autograd-compatible Mie solver, analytical Mie theory can be used as an error-free forward model through which the loss can be backpropagated to train the inverse network.
@@ IV.4 Autodiff multi-scattering: End-to-end metasurface design
As a final example, we demonstrate how PyMieDiff can be combined with the autodiff light scattering simulation toolkit “TorchGDM”,Ponomareva et al. (2025) to perform multi-particle scattering simulations. PyMieDiff evaluates the Mie response of one or several core-shell particles, which are then used by torchGDM to generate a structure model and calculate the optical response of an ensemble of many scatterers, where all optical interactions between the particles are taken into account. Since both tools fully support PyTorch automatic differentiation, gradient optimization can be used to iteratively optimize the positions and core/shell size parameters for any target observable. This is illustrated in figure 6a.
@@ V Limitations and perspectives
  • The recurrences currently require to be performed at double precision, which significantly slows down the code on consumer grade GPUs. In the future, more stable algorithms may be implemented such as logarithmic derivatives.Majic and Ru (2020); Zhang (2024)
  • Vector spherical harmonics are currently implemented only for order $l=1$, which limits the field calculations to particles with spherical symmetry. We plan to add pure pytorch implementations of general vector spherical harmonics that could be used for field calculations of general T-matrices. A possible route could be to expand on NVIDIA’s recent work about spherical Fourier neural operators.Bonev et al. (2023)
  • The current recurrences can become unstable in cases where many expansion orders are contributing to the response, for instance with very large particles or plasmonic/dielectric interfaces in larger particles. Be aware of these limitations. This could be mitigated in the future by implementing more stable algorithms, such as logarithmic derivatives recurrences.Bohren and Huffman (1998)
@@ VI Conclusions
At the time of the submission, we became aware of a manuscript developing similar ideas.Asadova et al. (2025a) This underpins the timeliness and importance of the differentiable formulation of algorithms that solve multiple-scattering problems, a key requirement for inverse design of photonic nanostructures.
@@ Acknowledgements.
For comparison and testing, we provide an alternative PyTorch-based scheme to compute the spherical Bessel functions $j_{n}$ using a continued fraction calculation,Lentz (1990) Using the Riccati–Bessel form $\psi_{n}(z)=z\,j_{n}(z)$ the three-term recurrence reads
@@
We truncate the continued fraction at a sufficiently large $n_{\max}$,Cai (2011) compute the $A_{n}$ by backward recurrence, reconstruct $\psi_{n}$ (and hence $j_{n}=\psi_{n}/z$) up to a multiplicative constant, and fix the global scale using $j_{0}(z)=\sin z/z$. Both implementations for $j_{n}$ are similarly stable and can be selected with the argument which_jn (“ratios” or “recurrence”).
@@
The vector spherical harmonics $\mathbf{N}_{lm}^{(k)}$ and $\mathbf{M}_{lm}^{(k)}$ can be used to describe the near fields of Mie scattering.Asadova et al. (2025b) For spherical symmetry, we need only the $l=1$ odd and even order harmonics, which are given by:Bohren and Huffman (1998)
@@
The Mie coefficients for a core-shell spherical particle are given byBohren and Huffman (1998)
@@
The Mie coefficients for internal fields can be found in literature.Bohren and Huffman (1998)
@@
where $\mu=\cos\theta$. The angular functions $\pi_{n}(\cos\theta)$ and $\tau_{n}(\cos\theta)$ are defined from the associated Legendre polynomials $P_{n}^{1}(\cos\theta)$.Bohren and Huffman (1998) These satisfy the recurrence relations
@@
The near-fields can be found by matching the continuity conditions at each spherical interface, which yields following formulae:Bohren and Huffman (1998)